OpenStudy (anonymous):

how do i get the derivative of y=x^4e^x

OpenStudy (anonymous):

hey, good morning, mimi

OpenStudy (anonymous):

use the product rule :) y' = u'v + uv' where : - u = x^4 - u' = 4x^3 - v = e^x - v' = e^x just plug them in ^_^

OpenStudy (anonymous):

good morning, when is your exam? what time?

OpenStudy (anonymous):

in 3 hours

OpenStudy (anonymous):

how do you feel?

OpenStudy (anonymous):

I am just chilling now :)

OpenStudy (anonymous):

great

OpenStudy (anonymous):

did you eat good?

OpenStudy (anonymous):

I got x^4e^x+4x^3e^x...is that right???

OpenStudy (anonymous):

sleep good?

OpenStudy (anonymous):

no, but I will in a moment, Mom

OpenStudy (anonymous):

yep roam :) that's correct ^_^

OpenStudy (anonymous):

>_> haha, very funny, I'm nervous lol

OpenStudy (anonymous):

well, what are you waiting for, go EAT, heavy breakfast go go go

OpenStudy (anonymous):

don't worry Mom, I will eat good, relax, get all my things ready, my ID, my maths instruments, I will be alright

OpenStudy (anonymous):

stop calling me mom lol, did you sleep well?

OpenStudy (anonymous):

=D alright, I am going to eat now, bye

OpenStudy (anonymous):

yes, slept well

OpenStudy (anonymous):

okay, good good

OpenStudy (anonymous):

wait a min, what time did you go to bed?

OpenStudy (anonymous):

so if i set it to zero, what would my critical numbers be?? sorry i am really bad at exponential and log functions

OpenStudy (anonymous):

you'll have to factor first to find the zeros, lol it's alright, when you factor and get the zeros, your zeros will be the critical numbers and to find the points , you'll have to plug them in the original function to get the y's :)

OpenStudy (anonymous):

ok thanks sstarica, you are really helping me out tonight! :)

OpenStudy (anonymous):

np :) good luck!

OpenStudy (anonymous):

Good luck today, oktal

OpenStudy (anonymous):

he went to eat lol

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i should be doing my work anyway

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

done?

OpenStudy (anonymous):

>_> eat well

OpenStudy (anonymous):

yeah..

OpenStudy (anonymous):

now I am all ready :)

OpenStudy (anonymous):

all I have to do is go to the university campus, find the place where the exams will take place and do my duty :)

OpenStudy (anonymous):

great ^_^ , don't panic if you don't know how to solve one question, if you take alot of time on it, skip it and proceed to the other questions, when you're done and there's still time, go back to the questions that you've skipped

OpenStudy (anonymous):

and you'll do just fine ^_^

OpenStudy (anonymous):

Good luuuuuuuuuccckkkkk

OpenStudy (anonymous):

lol , yep :)

OpenStudy (anonymous):

You'll be fine. Just be strategic. Don't spend too much time on a question if you're stuck.

OpenStudy (anonymous):

exactly~

OpenStudy (anonymous):

oh, keep a bottle of cold water near you

OpenStudy (anonymous):

I dont keep bottles of cold water near me... when I lived in Africa for some time, I used to because it is so freaking hot there and my body needs alot, but here, in this cold mystical country, I might prefare hot chocolate instead :)

OpenStudy (anonymous):

If I were you, I'll stay away from hot chocolate before any test, it has milk and this will make you drowsy

OpenStudy (anonymous):

Last semester, before my calculus I midterm exam in an hour, I drank hot chocolate assuming it to give me energy, but instead I got drowsy and sleepy and messed up the test

OpenStudy (anonymous):

nah.. I wont take anything to the exam.... because i will concentrate really hard on the questions and I will forget to drink anyway...

OpenStudy (anonymous):

you'll have breaks right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

If you can get your hands on Powerade or something, drink that

OpenStudy (anonymous):

how long is it? ._.

OpenStudy (anonymous):

the exam is VERY small and short

OpenStudy (anonymous):

lol, loki that's not good for the heart, not even redbull ._.

OpenStudy (anonymous):

powerade just replaces electolytes

OpenStudy (anonymous):

good, small short, easy simple, and you'll be done in an hour :)

OpenStudy (anonymous):

his nerves have to work

OpenStudy (anonymous):

his nerves are working just fine after such a heavy breakfast

OpenStudy (anonymous):

I don't like those power drinks

OpenStudy (anonymous):

not healthy

OpenStudy (anonymous):

it is only 3.5 hours

OpenStudy (anonymous):

take chocolate with you just in case

OpenStudy (anonymous):

and thats maths, some physics, some chem, and an engish essay and motivation letter

OpenStudy (anonymous):

too much concentration will cause you to get hungry

OpenStudy (anonymous):

especially when you have 3 sciences

OpenStudy (anonymous):

I guess Red Bull is a better choice, once I remember drinking like 4 cans, and it made me focus like a laser :)

OpenStudy (anonymous):

._. redbull is NOT GOOD

OpenStudy (anonymous):

and 4 cans? are you crazy?!

OpenStudy (anonymous):

you got luck in not getting a heart attack that's for sure.

OpenStudy (anonymous):

lucky*

OpenStudy (anonymous):

red bull's awesome

OpenStudy (anonymous):

who cares if it is good or not, what I care about is to concentrate and have an energetic mind

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

yeah, you'll care when you fall ill won't you?

OpenStudy (anonymous):

your heart repairs

OpenStudy (anonymous):

no it doesn't

OpenStudy (anonymous):

don't take it that slightly

OpenStudy (anonymous):

it does...i saw it on oprah once

OpenStudy (anonymous):

you have to take care of yourself , both of you.

OpenStudy (anonymous):

well oprah is not always true

OpenStudy (anonymous):

the doctor was

OpenStudy (anonymous):

all organs can slowly repaire themselves, yeah I saw it somewhere too... the guys name was Paul Bragg or something, he was this health pioneer

OpenStudy (anonymous):

I know someone who has a weak heart, that cannot be repaired

OpenStudy (anonymous):

that's genetic, though

OpenStudy (anonymous):

4 cans? what has gotten into you.

OpenStudy (anonymous):

his repair mechanisms are shot from birth

OpenStudy (anonymous):

not genetic loki

OpenStudy (anonymous):

it just happened

OpenStudy (anonymous):

my repair mechanisms are GREAT, I once got "messed up" and had some bones broken in the leg and fist and wrist... but in about 1 month I was able to perform just as I did before....

OpenStudy (anonymous):

effects won't show up now, but they will certainly will later , just be careful , try not to drink those power drinks too much guys.

OpenStudy (anonymous):

you are young, what about later?

OpenStudy (anonymous):

it'd be just one drink of red bull ... ;)

OpenStudy (anonymous):

it is not like I drink red bull like water every day.... that would waste too much money, so I can't :)

OpenStudy (anonymous):

well, try not to drink too much , both

OpenStudy (anonymous):

AND my dear, I don't drink any alcohol, so I have to "conpensate" with Red Bull :)

OpenStudy (anonymous):

yeah I know, again not too much

OpenStudy (anonymous):

^_^